OpenStudy (anonymous):

What is the correct solution to the equation? 8(x + 2) = 24 A. 1 B. 3 C. 5 D. 64

jigglypuff314 (jigglypuff314):

8(x + 2) = 24 distributive property 8x + 16 = 24 subtract 16 from both sides -16 -16 8x = 8 divide both sides by 8 /8 /8 x = 8 divided by 8 = (simplify)

OpenStudy (anonymous):

So the answer would be .. 1? Im kind of tired so lol

jigglypuff314 (jigglypuff314):

correct! :) lol

OpenStudy (anonymous):

Alright thanks!: )